Suppose a company has a forklift but is considering purchasing a new electric lift truck that would cost $3000000 and has operating cost of $50500 in the first year. For the remaining years, operating costs increase each year by 35% over the previous year's operating costs. It loses 7% of its value every year from the previous year's salvage value. The lift truck has a maximum life of 6 years. The firm's required rate of return is 5%. Find the economic service life of this new machine. 
Answer #1

Solution :

Year Operating Cost P/F (5%,n) Present value of Operating cost at the beginning of year Cumulative Present value of Operating cost at the beginning of year Present worth of cumulative operating cost A/P (5%,n) Average equivalent annual costs
1 2 3 4 = 2 * 3 5 6 = 5 + 3000000 7 8 = 6 * 7
1                         50,500 0.9524                               48,096                             48,096                        3,048,096 1.0500                                 3,200,501
2                         68,175 0.907                               61,835                           109,931                        3,109,931 0.5378                                 1,672,521
3                         92,036 0.8638                               79,501                           189,432                        3,189,432 0.3673                                 1,171,478
4                       124,249 0.8227                            102,220                           291,651                        3,291,651 0.2820                                     928,246
5                       167,736 0.7835                            131,421                           423,073                        3,423,073 0.1728                                     591,507 LEAST
6                       226,444 0.7462                            168,972                           592,045                        3,592,045 0.1971                                     707,992

Since the annual cost is least in 5 years, so, economic service life of the machine = 5 years

**** In case of any querry, Please comment below.

**** If this answer is helpful to you, Please give a thumbs up.